# Physics-informed DeepCT: Sinogram Wavelet Decomposition Meets Masked Diffusion

### v1

@@ 4.1.1 AAPM Challenge Data
The dataset used for training and testing is part of the AAPM Low-Dose Challenge[44] provided by the Mayo Clinic. It consists of 5,388 CT slices from 10 patients. In this study, 4,742 slices from 9 patients were used for training, while 12 slices from the remaining patient were used for testing. Each slice is 1 mm thick. To evaluate the performance of the proposed method with different sparse views, CT images were reconstructed using 30, 60, 90 and 120 projections, and both qualitative and quantitative analyses were performed. Reference images are generated from 720 projection views using the FBP algorithm. For the fan-beam CT reconstruction task, ray-driven algorithms[45, 46] were employed to simulate the generation of projection data. The distance from the center of rotation to the source and detector is set at 40 cm and 40 cm, respectively. The detector is 41.3 cm wide, has 720 detector elements, and a total of 720 projection views are evenly distributed.
4.1.2 CIRS Phantom Data
@@
We use a dataset of high-quality CT scans to further analyze the performance of the proposed method, each with dimensions of 512×512×100 voxels, and a voxel dimension of 0.78×0.78×0.625 mm3. The data set was collected using the GE Discovery HD750 CT scanner combined with a bionic model provided by CIRS. The vacuum tube current is set to 600 mAs, the source-wheelbase separation of the CT system is 573mm, and the source-detector distance is 1010 mm. To evaluate the efficacy of the reconstruction algorithms across different sparse-view sampling scenarios, CT images were reconstructed using 30, 60, 90 and 120 projection, the reconstruction results underwent both qualitative and quantitative analysis.
4.2 Implementation Details
@@
In the experimental part of this study, we use Adam algorithm to train the model, and the learning rate is set to $10^{-3}$. At the same time, use the Kaiming initialization to initialize the weights of the model. Three kinds of simulation masks are designed in this method. The first simulates the 10, 20, 30, 60, 90, 120, 180, and 720 projection angles of the sparse view mask. The second is three randomly positioned circle masks with a radius of 48 pixels; The third is a mask one-fifth of the width of the original image. At the same time, use the Kaiming initialization to initialize the weights of the model. The method is implemented in Python using Operator Discretization Library (ODL)[47] and PyTorch on a personal workstation with a GPU card (Tesla V100-PCIE-16 GB). In the reconstruction process, set the iteration parameter of $N<1200$. Our source code is publicly available on GitHub and can be accessed via the following link: https://github.com/yqx7150/SWARM. To evaluate the performance quantitatively, peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) and mean square error (MSE) were used.
4.3 Reconstruction Experiments
4.3.1 AAPM Reconstruction Results
@@
Table 1: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of AAPM Challenge Data Using Different Methods at 30, 60, 90, 120 Views.
Views FBP[48] FBPConvNet[14] HDNet[18] NCSN++[32] GMSD[21] SWARM
30 19.16/0.4596/12.561 29.76/0.9403/1.175 28.30/0.9347/2.070 32.66/0.9445/0.596 31.11/0.7888/0.849 34.38/0.9585/0.415
60 23.28/0.5957/4.815 34.23/0.9564/0.402 35.28/0.9706/0.345 36.97/0.9723/0.232 36.29/0.8445/0.276 38.65/0.9779/0.163
90 26.32/0.7088/2.388 36.25/0.9611/0.260 38.45/0.9809/0.164 39.64/0.9820/0.116 39.22/0.8745/0.127 41.76/0.9877/0.070
120 28.50/0.7961/1.447 38.03/0.9692/0.161 39.21/0.9851/0.151 40.82/0.9854/0.093 40.53/0.9060/0.097 42.96/0.9901/0.053
@@
Figure 3: Reconstruction images from 120 views using different methods with AAPM challenge data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) NCSN++, (f) GMSD and (g) SWARM. The display window is [0, 50] HU. The second line is the residual between the reference image and the reconstructed image.
Table 2: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of CIRS Phantom Data Using Different Methods at 30, 60, 90, 120 Views.
Views FBP[48] FBPConvNet[14] HDNet[18] NCSN++[32] GMSD[21] SWARM
30 12.67/0.4002/54.086 23.93/0.8850/4.052 27.27/0.9222/1.898 27.30/0.9287/1.923 27.54/0.7914/1.785 30.58/0.9618/0.899
60 17.15/0.5057/19.294 24.25/0.9308/3.762 32.53/0.9750/0.562 34.88/0.9765/0.329 33.07/0.8675/0.499 38.45/0.9886/0.144
90 21.78/0.6215/6.636 29.60/0.9320/1.102 37.21/0.9881/0.193 40.03/0.9879/0.101 37.75/0.8913/0.170 44.70/0.9957/0.035
120 25.08/0.6940/3.102 31.19/0.9444/0.764 40.20/0.9916/0.970 42.73/0.9920/0.054 39.86/0.9148/0.104 46.68/0.9968/0.022
@@
Figure 4: Reconstruction images from 60 views using different methods with CIRS phantom data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) NCSN++, (f) GMSD and (g) SWARM. The display window is [0,180] HU. The second line is the residual between the reference image and the reconstructed image.
@@
In order to assess the effectiveness of the proposed algorithm, we will evaluate the performance and compare some representative methods including FBP[48], FBPConvNet[14], HDNet[18], NCSN++[32], GMSD[21]. Sparse-view CT reconstructions were conducted using 30, 60, 90 and 120 projections, respectively. Table 1 presents the PSNR, SSIM, and MSE values for the reconstruction results of the AAPM challenge dataset. The best values for reconstructed images with different projection views are highlighted in bold. SWARM demonstrates superior performance over other reconstruction techniques in terms of mean values for the AAPM test data. In contrast, FBP and HDNet exhibit inferior numerical performance for CT reconstruction under sparse-view conditions. Furthermore, SWARM shows significant advantages over existing technologies.
Fig. 3 shows the visual reconstruction results of the AAPM test dataset. The traditional FBP reconstruction exhibits poor image quality, characterized by blurred details and significant edge artifacts. Although FBPConvNet has shown some improvements, it still suffers from edge blurring and a loss of finer details. The complex dual-domain iterative model HDNet, while effective in certain aspects, leads to information loss due to excessive smoothing. Additionally, NCSN++ and GMSD reconstruction also results in a noticeable loss of texture details and the presence of edge artifacts after iterative processing. In contrast, SWARM not only effectively eliminates edge artifacts but also preserves more structural and textural details, and the reconstruction results are closer to groundtruth.
4.3.2 CIRS Phantom Reconstruction Results
@@
To further evaluate the generalization and robustness of the proposed unsupervised learning method, prior knowledge was learned from the AAPM challenge dataset, and the model performance was assessed using CIRS simulated test data. Table 2 show that SWARM significantly outperforms the other approaches in terms of quantitative metrics. The image quality of FBP and FBPConvNet on the CIRS phantom dataset is notably inferior, indicating considerable loss of image details. Although the quantitative metrics of HDNet, NCSN++ and GMSD remain relatively stable, their reconstruction performance still shows a noticeable decline. In contrast, SWARM got the best results, demonstrating its strong generalization and robustness.
As shown in Fig. 4, our method achieves excellent visual results. The images reconstructed using FBP exhibit significant streak artifacts and detail loss. While FBPConvNet shows some improvement, the edges and textures remain blurred. HDNet enhances artifact removal and detail recovery, but the internal structure edges are still unclear. NCSN++ and GMSD effectively reduce noise, but excessive smoothing causes blurring of textures and structures. In contrast, our method not only effectively removes artifacts but also preserves rich texture details, resulting in better visual quality, which fully demonstrates its generalization capability and robustness.
@@
4.4 Profile Lines Analysis
In this study, the edge preserving performance of different reconstruction methods is evaluated by comparing the degree of matching between the generated contours and the real contours. As shown in Fig. 5, in the comprehensive comparison, the contour generated by SWARM is closer to the real contour, and the contour line is the most accurate. The contours of other methods have poor consistency with the real contours. In addition, in areas with significant gray scale changes, the consistency between the contours generated by SWARM method and the real data is more obvious, which proves its robustness in contour preservation.
Table 3: Reconstruction Results of AAPM Challenge Data
@@ 4.4 Profile Lines Analysis
SWARM 30 34.38 0.9585 0.415
60 38.65 0.9779 0.163
90 41.76 0.9877 0.070
120 42.96 0.9901 0.053
Figure 6: Reconstruction images obtained from 60 views using different methods. (a) The reference image versus the images reconstructed by (b) SHD, (c) SRM, (d) SWARM. The display window is [0, 230] HU. The second line is the residual between the reference image and the reconstructed image.
4.5 Ablation and Generalization Study
4.5.1 Different Components in SWARM Model
SWARM uses a dual-diffusion model for iterative reconstruction, where SRM is trained using a random mask strategy and SHD is trained based on the random high-frequency components of the sinogram. Quantitative analysis in Table 3 shows that SWARM significantly enhances reconstruction. Fig. 6 further illustrates the influence of the model combination, and the qualitative results demonstrate that the cascaded iterative reconstruction of SRM and SHD fully leverages their complementary advantages, significantly improving image reconstruction quality. This demonstrates the effectiveness of our approach, leveraging the strengths of both models to achieve superior reconstruction.
4.5.2 Effectiveness of Random Masking and High-Frequency of Sinogram
As shown in Table 4 and Fig. 7, we study the effects of random masks and randomness based on high-frequency components in the wavelet domain in terms of quantitative indexes and visual effects respectively. NMS is the non-mask for sinogram, and NRH is the non-random wavelet for sinogram. Generalization experiments on the CIRS test set further confirm that the embedded random mask not only greatly improves the generalization ability of the model, but also enhances the robustness of the model, thus verifying the effectiveness of the proposed method. In addition, the training strategy that randomly selects high-frequency components as the input of network training has good stability, effectively enhances the generalization ability of the model, and thus improves the overall reconstruction efficiency.
Table 4: Reconstruction Results of CIRS Phantom Data
@@ 4.5.2 Effectiveness of Random Masking and High-Frequency of Sinogram
SWARM 30 30.58 0.9618 0.899
60 38.45 0.9886 0.144
90 44.70 0.9957 0.035
120 46.68 0.9968 0.022
Figure 7: Reconstruction images obtained from 90 views using different methods. (a) The reference image versus the images reconstructed by (b) SHD+NMS, (c) SRM+NRH, (d) SWARM. The display window is [0, 120] HU. The second line is the residual between the reference image and the reconstructed image.
5 DISCUSSIONS AND CONCLUSION
In this study, we proposed a diffusion strategy with random masking and high-frequency wavelet decomposition of sinograms for SVCT reconstruction. This method significantly improved the performance of the model in terms of global consistency and detail reconstruction, thus enhancing the accuracy and reliability of the reconstructed image. Notably, tests conducted on various datasets have validated SWARM’s excellent performance in terms of generalization and robustness. However, the proposed method faced certain challenges. The inherent characteristics of the diffusion model led to longer computation times, and the conversion between the projection domain and the wavelet domain added additional computational cost. Future research could explore diffusion models in latent spaces to reduce computational complexity and enhance efficiency while maintaining high performance.
[…]

### v2

@@ IV-A Data Specification
1) AAPM Challenge Data: The dataset used for training and testing is part of the AAPM Low-Dose Challenge[49] provided by the Mayo Clinic. In this study, 4,742 slices from 9 patients were used for training, while 12 slices from the remaining patient were used for testing. Each slice is 1 mm thick. To evaluate the performance of the proposed method with different sparse views, CT images were reconstructed using 60, 90, and 120 projections, and both qualitative and quantitative analyses were performed. Reference images are generated from 720 projection views using the FBP algorithm. For the fan-beam CT reconstruction task, ray-driven algorithms[50, 51] were employed to simulate the generation of projection data. The distance from the center of rotation to the source and detector is set at 40 cm and 40 cm, respectively. The detector is 41.3 cm wide, has 720 detector elements, and a total of 720 projection views are evenly distributed.
@@
2) CIRS Phantom Data: We use a dataset of high-quality CT scans to further analyze the performance of the proposed method, each with dimensions of 512×512×100 voxels, and a voxel dimension of 0.78×0.78×0.625 mm3. The data set was collected using the GE Discovery HD750 CT scanner combined with a bionic model provided by CIRS. The vacuum tube current is set to 600 mAs, the source-wheelbase separation of the CT system is 573mm, and the source-detector distance is 1010 mm. To evaluate the efficacy of the reconstruction algorithms across different sparse-view sampling scenarios, CT images were reconstructed using 60, 90, and 120 projection, the reconstruction results underwent both qualitative and quantitative analysis.
3) Dental Arch Data: The data set is the clinical data collected by JIROX Dental CBCT device produced by YOFO (Hefei) Medical Technology Co., Ltd. The source-to-image distance (SID) of the device is 1700 mm, and the source-to-detector distance (SAD) is 1500 mm. The device operates at a voltage of 100 kV and a current of 6 mA. The detector array consisted of 768 × 768 elements, with each detector element measuring 0.2 × 0.2 mm. The dataset consists of 20 cases, and each case provides 200 slices. The sinograms have a size of 1536×1600, which correspond to 1600 uniformly-sampled projections within the angular range of $[0^{\circ},360^{\circ}]$. The dataset consists of full-view sinogram and their corresponding FBP reconstructions, with 512x512 image matrix. We randomly selected 1 case from the obtained image matrices and performed CT reconstruction using 60, 90, and 120 projections, respectively.
IV-B Implementation Details
@@
In the experimental part of this study, we use Adam algorithm to train the model, and the learning rate is set to $10^{-3}$. The virtual mask are randomly designed in the following way. By simulating the sparse-view mask at 10, 20, 30, 60, 90, 120, and 180 projection. At the same time, use the Kaiming initialization to initialize the weights of the model. The method is implemented in Python using Operator Discretization Library (ODL)[52] and PyTorch on a personal workstation with a GPU card (Tesla V100-PCIE-16 GB). During the reconstruction process, the iteration parameter was set to $N=800$ for 60 views and $N=1200$ for 90 and 120 views. Our source code is publicly available on GitHub and can be accessed via the following link: https://github.com/yqx7150/SWARM. To evaluate the performance quantitatively, peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) and mean square error (MSE) were used.
@@
TABLE I: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of AAPM Challenge Data Using Different Methods at 60, 90, 120 Views.
Views FBP[53] FBPConvNet[14] HDNet[18] GMSD[21] SWORD[22] SWARM
60 23.28/0.5957/4.815 34.23/0.9564/0.402 35.28/0.9706/0.345 36.29/0.9684/0.276 37.09/0.9738/0.213 38.43/0.9809/0.188
90 26.32/0.7088/2.388 36.25/0.9611/0.260 38.45/0.9809/0.164 39.22/0.9800/0.127 40.14/0.9835/0.103 41.91/0.9882/0.067
120 28.50/0.7961/1.447 38.03/0.9692/0.161 39.21/0.9851/0.151 40.53/0.9846/0.097 42.04/0.9884/0.066 42.80/0.9904/0.058
@@
Figure 3: Reconstruction images from 120 views using different methods with AAPM challenge data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) GMSD, (f) SWORD, and (g) SWARM. The display window is [-50, 490] HU. The second line is the residual between the reference image and the reconstructed image.
@@
Figure 4: Reconstruction images from 90 views using different methods with AAPM challenge data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) GMSD, (f) SWORD, and (g) SWARM. The display window is [-480, 945] HU. The second line is the residual between the reference image and the reconstructed image.
TABLE II: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of CIRS Phantom Data Using Different Methods at 60, 90, 120 Views.
Views FBP[53] FBPConvNet[14] HDNet[18] GMSD[21] SWORD[22] SWARM
60 17.15/0.5057/19.294 24.25/0.9308/3.762 32.53/0.9750/0.562 33.07/0.9744/0.499 32.89/0.9750/0.520 38.38/0.9887/0.146
90 21.78/0.6215/6.636 29.60/0.9320/1.102 37.21/0.9881/0.193 37.75/0.9866/0.170 39.21/0.9901/0.124 44.73/0.9958/0.034
120 25.08/0.6940/3.102 31.19/0.9444/0.764 40.20/0.9916/0.970 39.86/0.9915/0.104 41.30/0.9935/0.074 46.80/0.9969/0.021
Figure 5: Reconstruction images from 60 views using different methods with CIRS phantom data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) GMSD, (f) SWORD, and (g) SWARM. The display window is [675, 1300] HU. The second line is the residual between the reference image and the reconstructed image.
TABLE III: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of Dental Arch Data Using Different Methods at 60, 90, 120 Views.
Views FBP[53] FBPConvNet[14] HDNet[18] GMSD[21] SWORD[22] SWARM
60 25.55/0.7015/2.818 31.86/0.9492/0.678 31.20/0.9658/0.972 33.45/0.9678/0.483 36.33/0.9757/0.251 40.22/0.9911/0.099
90 28.12/0.7917/1.574 34.04/0.9574/0.406 36.55/0.9232/0.248 37.70/0.9826/0.183 39.93/0.9890/0.109 44.63/0.9955/0.037
120 29.87/0.8519/1.056 35.55/0.9716/0.307 40.25/0.9904/0.102 40.68/0.9902/0.091 43.58/0.9944/0.047 47.11/0.9971/0.022
Figure 6: Reconstruction images from 60 views using different methods with Dental Arch data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) GMSD, (f) SWORD, and (g) SWARM. The display window is [-60, 1300] HU. The second line is the residual between the reference image and the reconstructed image.
IV-C Reconstruction Experiments
@@
1) AAPM Reconstruction Results: In order to assess the effectiveness of the proposed algorithm, we evaluated the performance and compared some representative methods including FBP[53], FBPConvNet[14], HDNet[18], GMSD[21], SWORD[22]. Sparse-view CT reconstructions were conducted using 60, 90, and 120 projection, respectively. Table I presents the PSNR, SSIM and MSE values for the reconstruction results of the AAPM challenge dataset. The best values for reconstructed images with different projection views are highlighted in bold. SWARM demonstrated superior performance over other reconstruction techniques in terms of mean values for the AAPM test data. In contrast, FBP and HDNet exhibited inferior numerical performance for CT reconstruction under sparse-view conditions. Furthermore, SWARM showed significant advantages over existing technologies.
Fig. 3 and Fig. 4 shows the visual reconstruction results of the AAPM test dataset. FBP exhibit suboptimal performance in image reconstruction, often resulting in severe blurring of fine details and pronounced edge artifacts. Although FBPConvNet has alleviated some of these limitations, it still suffers from insufficient edge sharpness and inadequate recovery of fine structures, leading to the omission of subtle image features. The dual-domain iterative model HDNet, which integrates information from both frequency and image domains, achieves certain improvements in reconstruction quality. However, it tends to introduce excessive smoothing, consequently causing the loss of critical structural information. Similarly, GMSD and SWORD inevitably compromise texture information during iterative reconstruction, thereby diminishing the realism and fidelity of reconstructed images. In contrast, the SWARM method demonstrates superior performance in preserving texture details and restoring edge contours. It maintains higher fidelity in representing structural features such as the shape and grayscale distribution of target regions, exhibiting significantly improved consistency with ground truth images. Both in terms of visual perception and quantitative metrics, SWARM consistently outperforms the aforementioned approaches.
@@
2) CIRS Phantom Reconstruction Results: To further evaluate the generalization and robustness of the proposed unsupervised learning method, prior knowledge was learned from the AAPM challenge dataset, and the model performance was assessed using CIRS phantom data. Table II shows that SWARM significantly outperforms the other approaches in terms of quantitative metrics. The evaluation results of FBP and FBPConvNet on the CIRS phantom dataset are notably inferior, indicating considerable loss of image information. Although the quantitative metrics of HDNet and GMSD remain relatively stable, their reconstruction performance still shows a noticeable decline. SWORD shows relatively stable reconstruction performance in various indicators. However, it can still be observed that its reconstruction quality decreases to a certain extent in some test cases, revealing the limitation of insufficient generalization ability. In contrast, SWARM gets the best results, demonstrating its strong generalization and robustness.
As shown in Fig. 5, our method achieves excellent visual results. Images reconstructed using the conventional FBP method suffer from prominent streak artifacts and significant loss of structural details. FBPConvNet remains limited in restoring edge sharpness and texture details, resulting in blurred structural contours. Despite leveraging iterative optimization of dual-domain information, HDNet remains inadequate in accurately delineating internal structural boundaries. Similarly, GMSD and SWORD tend to over-smooth the reconstructed images, leading to the loss of texture information and blurring of fine structures. In contrast, SWARM demonstrates superior performance by effectively suppressing artifacts while preserving rich texture details and structural features. This results in a substantial enhancement of visual reconstruction quality, fully showcasing the method’s excellent generalization capability and robustness across diverse reconstruction scenarios.
3) Dental Arch Reconstruction Results: To evaluate the practicality of the proposed method, the model learned prior knowledge from the abdominal dataset of the AAPM challenge and assessed the model performance using the Panoramic Dental Arch data. Table III shows the significant advantages of SWARM in terms of quantitative metrics. Although other comparison methods remain relatively stable in the evaluation results or show slight improvements, they still fail to fully demonstrate their reconstruction capabilities. In contrast, SWARM shows obvious advantages with strong generalization ability and robustness, which provides an effective idea for solving the problem of SVCT reconstruction, and is expected to provide a potential reference for clinical diagnosis in stomatology in theory.
As shown in Fig. 6 and Fig. 7, our approach provides excellent visual results in the processing of oral data. In contrast, images reconstructed using FBP, FBPConvNet, HDNet, and GMSD exhibit significant streak artifacts and detail loss, particularly in the dental regions. While SWORD alleviates some of these issues, it still struggles with the restoration of structural details, especially in areas such as the mandibular molars, resulting in blurred images. In comparison, SWARM not only effectively suppresses streak artifacts but also achieves notable success in preserving texture details, particularly in the high-quality structural reconstruction of the mandibular molars. Overall, SWARM stands out in terms of visual reconstruction quality for oral images, providing clinicians with clearer and more accurate imaging support, thus enhancing the reliability and precision of diagnostic assistance.
Figure 7: Reconstruction images from 90 views using different methods with Dental Arch data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) GMSD, (f) SWORD, and (g) SWARM. The display window is [-500, 970] HU. The second line is the residual between the reference image and the reconstructed image.
@@
IV-D Profile Lines Analysis
In this study, we evaluated the edge-preserving performance of different reconstruction methods by comparing the generated contours with the ground truth contours. As shown in Fig. 8, the contours generated by FBP, FBPConvNet, HDNet, and SWORD exhibit poor alignment with the ground truth, with noticeable deviations in edge details. In contrast, the contours generated by GMSD and SWARM exhibit a closer match to the ground truth, further highlighting the advantage of diffusion models in preserving edge details. Among these methods, the contours generated by SWARM are the most accurate, particularly in regions with significant gray-scale variations, where the consistency between SWARM generated contours and the actual data is most evident. These results further validate the superiority and robustness of the SWARM method in edge preservation.
Figure 9: Boxplot of Performance on CIRS Test Dataset Under Different Mask Rates (0%, 10%, 20%, 30%, 50%, 70%).
Figure 10: Reconstruction images obtained from 60 views using different methods. (a) The reference image versus the images reconstructed by (b) SHD, (c) SRM, (d) SWARM. The display window is [-180, 1300] HU. The second line is the residual between the reference image and the reconstructed image.
IV-E Ablation and Generalization Study
1) Mask Ratio Analysis: To systematically evaluate the impact of different mask sparsity ratios on image reconstruction performance, ablation experiments were conducted with mask ratios of 0%, 10%, 20%, 30%, 50%, and 70%, as illustrated in Fig. 9. As the masking ratio increases, both PSNR and SSIM initially improve, plateau, and then decline, while MSE demonstrates a typical “U-shaped” curve-decreasing initially and rising again beyond a certain point. This trend suggests that extremely low or high masking ratios hinder the model’s ability to effectively infer missing information, resulting in suboptimal reconstruction quality. In contrast, a moderate masking level (e.g., 20%) introduces a beneficial degree of sparsity that stimulates the model’s prior learning capacity, achieving an optimal balance between reconstruction quality and robustness. These findings further validate the hypothesis proposed in the methodological motivation: introducing an appropriate level of random masking enhances the model’s ability to complete unobserved regions, thereby improving overall reconstruction performance. This provides clear experimental evidence and guidance for selecting optimal mask ratios in future applications.
2) Different Components in SWARM: SWARM employs a dual-diffusion model for iterative reconstruction. During this process, SRM is trained using a random mask strategy, while SHD is trained based on random high-frequency components from the sinogram. The quantitative analysis in Table IV demonstrates that SWARM significantly improves reconstruction quality. Fig. 10 further illustrates the impact of the model combination. Qualitative results show that the cascade of SRM and SHD in iterative reconstruction effectively leverages their complementary advantages, leading to a significant enhancement in image reconstruction quality. This outcome validates the effectiveness of our approach, showing that combining the strengths of both models achieves superior reconstruction performance.
3) Effectiveness of Random Masking and High-Frequency of Sinogram: As shown in Table V and Fig. 11, we studied the effects of the random masking strategy and randomness based on high-frequency components in the wavelet domain from both quantitative metrics and visual perspectives. NMS represents the non-masked condition for the sinogram, while NRH represents the non-random wavelet condition for the sinogram. Generalization experiments conducted on the CIRS phantom data further confirmed that the incorporation of random masking not only significantly improved the model’s generalization ability but also enhanced its robustness, thereby validating the effectiveness of the proposed method. Additionally, the strategy of randomly selecting high-frequency components as network training inputs demonstrated good stability, effectively boosting the model’s generalization capability and improving overall reconstruction efficiency.
TABLE IV: Ablation Study of Two-Stage Model on AAPM Challenge Data.
@@ IV-E Ablation and Generalization Study
SWARM 60 38.43 0.9809 0.188
90 41.91 0.9882 0.067
120 42.80 0.9904 0.058
TABLE V: Generalization Study of Random Masking & Wavelet HF.
@@
SWARM 60 38.38 0.9887 0.146
90 44.73 0.9958 0.034
120 46.80 0.9969 0.021
IV-F Computational Complexity
As shown in Table VI, the comparison of the computational performance of various methods under 60 views, FBPConvNet demonstrates the optimal computational efficiency. However, FBPConvNet relies on separate training and solving for each scanning view. This angular dependency significantly increases the complexity of the sampling process and the training cost. In particular, it limits the practicality of the method in multi-view or dynamic imaging tasks. In contrast, although methods such as GMSD, SWORD, and SWARM have relatively higher time complexity, they only require training a single model for multi-angle reconstruction tasks. This significantly enhances the flexibility and generalization ability of these methods in application scenarios, especially in imaging environments with frequent angular changes or limited sampling conditions.
Figure 11: Reconstruction images obtained from 90 views using different methods. (a) The reference image versus the images reconstructed by (b) SHD+NMS, (c) SRM+NRH, (d) SWARM. The display window is [-180, 1370] HU. The second line is the residual between the reference image and the reconstructed image.
TABLE VI: Quantitative Evaluation of Reconstruction Time.
Methods FBPConvNet HDNet GMSD SWORD SWARM
Time 0.15s 1.27s 461.58s 1214.21s 1290.43s
V Discussion and Conclusion
In this study, we introduce a diffusion strategy for sparse-view computed tomography (SVCT) reconstruction that integrates random masking and high-frequency wavelet decomposition. The proposed method significantly enhances model performance in terms of global consistency and detail reconstruction, thereby improving the accuracy and reliability of reconstructed images. Testing across multiple datasets demonstrates that the SWARM approach exhibits strong generalization and robustness. However, the method faces challenges stemming from the inherent nature of diffusion models, which result in lengthy computation times for full sampling steps. Future research will explore optimizing the diffusion process via skip sampling to reduce computational complexity and enhance sampling efficiency while maintaining high model performance. Additionally, we plan to investigate mask designs compatible with CT scanning geometry and dynamically adaptive masks that incorporate anatomical features to further enhance the practicality and robustness of this innovative approach.